# A Personalised 3D+t Mesh Generative Model for Unveiling Normal Heart Dynamics

### v1

@@ Methods
Figure 1(a) illustrates the architecture of the proposed generative model, MeshHeart. Given a set of clinical conditions $c$, our goal is to develop a model that can generate a dynamic 3D cardiac mesh sequence, $X_{0:T-1}=\{x,x_{1},\cdots,x_{T-1}\}$, where $T$ denotes the number of time frames, that corresponds to the conditions $c$. Figure 1(b) shows an example of the input conditions and the generated mesh sequence. Without losing generality, we take age, sex, weight and height as conditions $c$ in this work. Age, weight and height are continuous variables, whereas sex is a binary variable. Each cardiac mesh $x_{t}=(v_{t},e_{t})$ is a graph with a set of vertices $v$ and a set of edges $e$ connecting them. The proposed generative model consists of several components, a mesh encoder $M_{enc}$ and a condition encoder $C_{enc}$ to project the mesh sequence and conditions into latent representations, a Transformer encoder $T_{enc}$ to capture the temporal dependency of the meshes at different time frames, and finally a Transformer decoder $T_{dec}$ and a mesh decoder $M_{dec}$ to decode the latent representations back to the mesh sequence. The mesh encoder $M_{enc}$, constructed using a graph convolutional network (GCN), extracts features from the cardiac mesh sequence $x_{0:T-1}$ and generates latent vectors $z_{0:T-1}$. The condition encoder $C_{enc}$, implemented with a multi-layer perceptron (MLP), maps the clinical conditions $c$ into a condition latent vector $z_{c}$ in the condition latent space. The mesh latent vectors $z_{0:T-1}$ at each time frame are concatenated with the condition latent vector $z_{c}$, which forms a sequence of input tokens to the Transformer decoder $T_{dnc}$. For simplicity, the concatenated vectors are still denoted as $z_{0:T-1}$. These components work together to learn the probability distribution of the cardiac mesh sequence $x$ conditioned on $c$, assuming that the latent space follows a Gaussian distribution. We seek a model $p_{\theta}(x|z_{c})$, parameterised by $\theta$, that is sufficiently flexible to learn to describe the mesh sequence $x$.
We assume a prior distribution $p(z_{a})$ over the latent variable $z_{a}$. The prior $p(z_{a})$ together with the decoder (constructed by $T_{dec}$ and $M_{dec}$) $p(x,z_{a}|z_{c})$, define a joint distribution, denoted as $p(x|z_{c})$. To train the model and perform inference, we need to compute the posterior distribution $p(z_{a}|x,z_{c})$, which is generally intractable. To turn the intractable posterior inference problem $p(z_{a}|x,z_{c})$ into a tractable problem, we introduce a parametric encoder model (constructed by $C_{enc},M_{enc},T_{enc}$) $q_{\phi}(z_{a}|x,z_{c})$ with $\phi$ to be the variational parameters, which approximates the true but intractable posterior distribution $p(z_{a}|x,z_{c})$ of the generative model, given an input $x$ and conditions $c$:
$q_{\phi}(z_{a}|x,z_{c})\approx p_{\theta}(z_{a}|x,z_{c})$ (1)
where $q_{\phi}(z_{a}|x,z_{c})$ often adopts a simpler form, e.g. the Gaussian distribution. By introducing the approximate posterior $q_{\phi}(z_{a}|x,z_{c})$, the log-likelihood of the conditional distribution $p_{\theta}(x|z_{c})$ for input data $x$, also known as evidence, can be formulated as:
$=\mathbb{E}_{z_{a}\sim q_{\phi}(z_{a}|x,z_{c})}\log\left[{\frac{p_{\theta}(x,z_{a}|z_{c})}{q_{\phi}(z_{a}|x,z_{c})}}\right]$
$+\mathbb{E}_{z_{a}\sim q_{\phi}(z_{a}|x,z_{c})}\log\left[\frac{q_{\phi}(z_{a}|x,z_{c})}{p_{\theta}(z_{a}|x,z_{c})}\right]$
where the second term denotes the Kullback-Leibler (KL) divergence $D_{KL}(q_{\phi}\parallel p_{\theta})$, between $q_{\phi}(z_{a}|x,z_{c})$ and $p_{\theta}(z_{a}|x,z_{c})$. It is non-negative and zero only if the approximate posterior $q_{\phi}(z_{a}|x,z_{c})$ equals the true posterior distribution $p_{\theta}(z_{a}|x,z_{c})$. Due to the non-negativity of the KL divergence, the first term in Eq. 2 is the lower bound of the evidence $\log[p_{\theta}(x|z_{c})]$, known as the evidence lower bound (ELBO). Instead of optimising the evidence $\log[p_{\theta}(x|z_{c})]$ which is often intractable, we optimise the ELBO:
$\max_{\theta,\phi}ELBO=\log[p_{\theta}(x|z_{c})]-D_{KL}$ (3)
@@
The Transformer encoder $T_{enc}$ comprises $L$ layers of alternating blocks of multi-head self-attention (MSA) and MLP. To ensure stability and effective learning, LayerNorm (LN) is applied before each block and residual connections are applied after each block. Similar to the additional class-token of vision Transformer [45], we append the input tokens $z_{0:T-1}$ with two learnable parameters $\mu_{token}$ and $\Sigma_{token}$, named as distribution parameter tokens. In the output layer, we extract the first two outputs $z^{L_{0}},z^{L_{1}}$ of the Transformer encoder as distribution parameters $\mu$ and $\Sigma$. We then use the reparameterisation trick [46] to derive a latent vector $z_{a}$ from $\mu$ and $\Sigma$, as shown in Figure 1(a). The encoding process is formulated as,
@@
$z_{a}=\mu+\epsilon\cdot\Sigma,\mu=z^{L_{0}},\Sigma=z^{L_{1}},\epsilon\sim\mathcal{N}(0,\mathbf{1})$
The learnt latent vector $z_{a}$ encodes the distribution information for the mesh sequence. It is concatenated with the condition latent vector $z_{c}$ and is provided as input to the Transformer decoder $T_{dec}$. A sequence of temporal tokens is fed to the Transformer decoder $T_{dec}$ as queries, providing temporal information for $T$ time frames. Inspired by the positional embedding in the vision Transformer [45], the temporal token $p_{t}$ at time frame $t$ is defined using the sinusoidal function with the same dimension $d$ as the latent vector $z_{a}$:
@@
where $i$ is the dimension index. The Transformer decoder generates a sequence of latent vectors, which are fed to the mesh decoder $M_{dec}$ composed of fully connected layers to reconstruct a 3D+t mesh sequence $X^{\prime}_{0:T-1}$.
@@
The training loss combines the mesh reconstruction loss $\mathcal{L}_{R}$, the Kullback-Leibler (KL) loss $\mathcal{L}_{KL}$, and a mesh smoothing loss $\mathcal{L}_{S}$. The mesh reconstruction loss $\mathcal{L}_{R}$ is defined as the Chamfer distance between the reconstructed mesh sequence $X^{\prime}_{0:T-1}=(V^{\prime},E^{\prime})$ and the ground truth $X_{0:T-1}=(V,E)$, formulated as $\mathcal{L}_{R}=\frac{1}{T}\sum_{t=0}^{T-1}D_{cham}(V^{\prime}_{t},V_{t})$, where $D_{cham}$ denotes the Chamber distance [47], $V^{\prime}_{t}$ and $V_{t}$ denote the mesh vertex coordinates for the reconstruction and the ground truth respectively:
@@
As in a VAE, the distribution of the latent space $\phi=\mathcal{N}(\mu,\Sigma)$ is encouraged to be close to a prior Gaussian distribution $\psi=\mathcal{N}(0,\mathbf{I})$. The KL divergence is defined between the latent distribution and the Gaussian prior distribution. To control the trade-off between distribution fitting and diversity, we adopt the $\beta$-VAE formulation [48]. The KL loss $\mathcal{L}_{KL}$ is formulated as
$\mathcal{L}_{KL}=\beta\cdot KL(\phi||\psi)$ (7)
which encourages the latent space $\phi=\mathcal{N}(\mu,\Sigma)$ to be close to the prior Gaussian distribution $\psi=\mathcal{N}(0,\mathbf{I})$.
@@
In terms of implementation, the mesh encoder $M_{enc}$ has three GCN layers and one fully connected (FC) layer. The mesh decoder $M_{dec}$ is composed of five FC layers. The Transformer encoder $T_{enc}$ and the decoder $T_{dec}$ consist of two layers, four attention heads, a feed-forward size of 1,024, and a dropout rate of 0.1. The latent vector dimensions for the mesh and the condition were set to 64 and 32. The weights $\beta$ and $\lambda_{s}$ in the loss function were empirically set to 0.01 and 1. The model was trained using the Adam optimiser with a fixed learning rate of $10^{-4}$ for $300$ epochs and with a batch size of one cardiac mesh sequence containing 50 time frames.
@@
Given conditions $c$, we draw 100 samples of the latent variable $z_{s}$ from a standard Gaussian distribution, specifically $z_{s}\sim\mathcal{N}(0,\mathbf{I})$, where $\mathbf{I}$ is the identity matrix. This sampling is intended to approximate the true underlying distribution of the data. The relationship between the latent variables and the observed data can be expressed as:
[…]
This equation indicates that by integrating over the latent variable $z_{s}$, we obtain the marginal probability of the observed data $x$ conditioned on $z_{c}$. Additionally, the joint probability of $x$ and $z_{s}$ can be decomposed as:
$p(x,z_{s})=p(x\mid z_{s},z_{c})p(z_{s}\mid z_{c})$ (11)
With MeshHeart, each cardiac mesh sequence $x$ can be projected into a latent vector $z$. We then define the latent vector $z$ to be last layer of the Transformer encoder $T_{enc}$. In this latent space, we define the latent delta $\Delta z$ as the difference between the latent vector of the real heart and the latent vector of the synthetic normal heart, using the Euclidean distance as the metric, formulated as,
$\Delta z=\sqrt{\sum_{i=1}^{n}(z^{\text{syth}}-z^{\text{real}})^{2}}$ (12)
The sequential cardiac meshes generated for both synthetic and real samples were encoded using the mesh encoder $M_{enc}$ and Transformer encoder $T_{enc}$. This process resulted in latent vectors $z^{\text{syth}}$ for synthetic samples and $z^{\text{real}}$ for real samples.
@@
This study used a data set of 38,309 participants obtained from the UK Biobank [30]. Each participant underwent cine cardiac magnetic resonance (CMR) imaging scans. From the cine CMR images, a 3D mesh sequence is derived to describe the shape and motion of the heart. The mesh sequence covers three anatomical structures, LV, Myo and RV. Each sequence contains 50 time frames over the course of a cardiac cycle. To derive cardiac meshes from the CMR images, automated segmentation [51] was applied to the image at the ED frame. A 3D template mesh [52] was fitted to the segmentation of the ED, resulting in an ED cardiac mesh. Subsequently, motion tracking was performed to warp the ED cardiac mesh across the time frames and obtain the entire mesh sequence. Motion tracking was performed using Deepali[53], which is the GPU-accelerated version of the non-rigid registration toolbox MIRTK[54]. All cardiac meshes maintained the same geometric structure, consisting of 22,043 vertices and 43,840 faces.
The dataset is partitioned into training/validation/test sets for developing the MeshHeart model and a clinical analysis set for evaluating its performance for disease classification task. Briefly, MeshHeart was trained on 15,000 healthy subjects from the Cheadle imaging centre. For parameter tuning and performance evaluation, MeshHeart was evaluated on a validation set of 2,000 and a test set of 4,000 healthy subjects, from three different sites, Cheadle, Reading, and Newcastle centres. For clinical analysis, including performing the disease classification study and latent delta PheWAS, we used a separate set of 17,309 subjects from the three imaging centres, including 7,178 healthy subjects and 10,131 subjects with cardiac diseases and hypertension. PheWAS was undertaken using the PheWAS R package with clinical outcomes and coded phenotypes converted to 1,163 categorical PheCodes. P-values were deemed significant with Bonferroni adjustment for the number of PheCodes. The details of the dataset split and the definition of disease code are described in Supplementary Table S1.
@@
The code for this research is available in the supplementary materials supplied as a zip file at the time of submission. It will be made publicly accessible on GitHub (https://github.com/MengyunQ/MeshHeart) before publication. The access information and links to the data repository will be updated once the publication is finalised.

### v2

@@ Methods
Figure 1(a) illustrates the architecture of the proposed generative model, MeshHeart. Given a set of clinical conditions $c$, our goal is to develop a model that can generate a dynamic 3D cardiac mesh sequence, $X_{0:T-1}=\{x_{0},x_{1},\cdots,x_{T-1}\}$, where $T$ denotes the number of time frames, that corresponds to the conditions $c$. Figure 1(b) shows an example of the input conditions and the generated mesh sequence. Without losing generality, we take age, sex, weight and height as conditions $c$ in this work. Age, weight and height are continuous variables, whereas sex is a binary variable. Each cardiac mesh $x_{t}=(v_{t},e_{t})$ is a graph with a set of vertices $v$ and a set of edges $e$ connecting them.
The proposed generative model consists of a mesh encoder $M_{\text{enc}}$, a Transformer encoder $T_{\text{enc}}$, a condition encoder $C_{\text{enc}}$, a Transformer decoder $T_{\text{dec}}$, and a mesh decoder $M_{\text{dec}}$. These components are designed to work together to learn the probability distribution $p_{\theta}(x|z_{c})$ of the cardiac mesh sequence conditioned on clinical attributes, where $\theta$ represents the decoder parameters, and $z_{c}$ denotes the condition latent vector. The condition encoder $C_{\text{enc}}$, implemented as a multi-layer perceptron (MLP), maps the clinical conditions $c$ into a condition latent vector $z_{c}$.
The mesh encoder $M_{\text{enc}}$, implemented as a graph convolutional network (GCN), processes the input cardiac mesh sequence $x_{0:T-1}$. It extracts latent representations $z_{0:T-1}$, where each vector $z_{t}$ corresponds to a latent representation of the cardiac mesh at time frame $t$. These latent vectors serve as intermediate representations of the cardiac mesh sequence.
@@
The latent vectors $z_{0:T-1}$ from the mesh encoder are concatenated with the condition latent vector $z_{c}$ to form a sequence of input tokens to the Transformer encoder $T_{\text{enc}}$. The Transformer encoder $T_{\text{enc}}$ captures temporal dependencies across the sequence, which comprises $L$ layers of alternating blocks of multi-head self-attention (MSA) and MLP. To ensure stability and effective learning, LayerNorm (LN) is applied before each block and residual connections are applied after each block. Similar to the class token in the vision Transformer [62], we append the input tokens $z_{0:T-1}$ with two learnable parameters $\mu_{token}$ and $\Sigma_{token}$, named as distribution parameter tokens, which parameterize a Gaussian distribution over the latent space. In the Transformer output layer, we extract the outputs from the distribution parameter tokens as distribution parameters $\mu$ and $\Sigma$. We then use the reparameterisation trick [63] to derive the latent $z_{a}$ from $\mu$ and $\Sigma$, as shown in Figure 1(a). The encoding process is formulated as,
@@
$z_{a}=\mu+\epsilon\cdot\Sigma,\epsilon\sim\mathcal{N}(0,\mathbf{1})$
The resulting latent vector $z_{a}$, derived after the reparameterisation step, captures the information about the distribution of the mesh sequence. This vector is concatenated with the condition latent vector $z_{c}$ to form the input to the Transformer decoder $T_{\text{dec}}$. The decoder uses these concatenated vectors as keys and values in the self-attention layer, while sinusoidal temporal positional encodings [62] serve as queries to incorporate temporal information. The temporal positional encoding $p_{t}$ at time frame $t$ is defined using the sinusoidal function with the same dimension $d$ as $z_{a}$:
@@
where $i$ denotes the dimension index. The Transformer decoder outputs a sequence of latent vectors, each corresponding to a mesh representation at a time point of the cardiac cycle. The latent vectors generated by the Transformer decoder are passed through the mesh decoder $M_{\text{dec}}$, composed of fully connected layers, to reconstruct the 3D+t cardiac mesh sequence $X^{\prime}_{0:T-1}$.
Probabilistic modelling and optimisation
Following the VAE formulation[63, 64], we assume a prior distribution $p(z_{a})$ over the latent variable $z_{a}$. The prior $p(z_{a})$, together with the decoder (constructed by $T_{\text{dec}}$ and $M_{\text{dec}}$), defines the joint distribution $p(x,z_{a}|z_{c})$. To train the model and perform inference, we need to compute the posterior distribution $p(z_{a}|x,z_{c})$, which is generally intractable. To turn the intractable posterior inference problem $p(z_{a}|x,z_{c})$ into a tractable problem, we introduce a parametric encoder model (constructed by $C_{\text{enc}},M_{\text{enc}},T_{\text{enc}}$) $q_{\phi}(z_{a}|x,z_{c})$ with $\phi$ to be the variational parameters, which approximates the true but intractable posterior distribution $p(z_{a}|x,z_{c})$ of the generative model, given an input $x$ and conditions $c$:
$q_{\phi}(z_{a}|x,z_{c})\approx p_{\theta}(z_{a}|x,z_{c})$ (3)
where $q_{\phi}(z_{a}|x,z_{c})$ often adopts a simpler form, e.g. the Gaussian distribution [63, 64]. By introducing the approximate posterior $q_{\phi}(z_{a}|x,z_{c})$, the log-likelihood of the conditional distribution $p_{\theta}(x|z_{c})$ for input data $x$, also known as evidence, can be formulated as:
$=\mathbb{E}_{z_{a}\sim q_{\phi}(z_{a}|x,z_{c})}\log\left[{\frac{p_{\theta}(x,z_{a}|z_{c})}{q_{\phi}(z_{a}|x,z_{c})}}\right]+\mathbb{E}_{z_{a}\sim q_{\phi}(z_{a}|x,z_{c})}\log\left[\frac{q_{\phi}(z_{a}|x,z_{c})}{p_{\theta}(z_{a}|x,z_{c})}\right]$
where the second term denotes the Kullback-Leibler (KL) divergence $D_{KL}(q_{\phi}\parallel p_{\theta})$, between $q_{\phi}(z_{a}|x,z_{c})$ and $p_{\theta}(z_{a}|x,z_{c})$ [63, 64]. It is non-negative and zero only if the approximate posterior $q_{\phi}(z_{a}|x,z_{c})$ equals the true posterior distribution $p_{\theta}(z_{a}|x,z_{c})$. Due to the non-negativity of the KL divergence, the first term in Eq. 4 is the lower bound of the evidence $\log[p_{\theta}(x|z_{c})]$, known as the evidence lower bound (ELBO). Instead of optimising the evidence $\log[p_{\theta}(x|z_{c})]$ which is often intractable, we optimise the ELBO:
$\min_{\theta,\phi}ELBO=-\log[p_{\theta}(x|z_{c})]+D_{KL}$ (5)
@@
Based on the ELBO, we define the concrete training loss function, which combines the mesh reconstruction loss $\mathcal{L}_{R}$, the Kullback-Leibler (KL) loss $\mathcal{L}_{KL}$, and a mesh smoothing loss $\mathcal{L}_{S}$. The mesh reconstruction loss $\mathcal{L}_{R}$ is defined as the Chamfer distance between the reconstructed mesh sequence $X^{\prime}_{0:T-1}=(V^{\prime},E^{\prime})$ and the ground truth $X_{0:T-1}=(V,E)$, formulated as $\mathcal{L}_{R}=\frac{1}{T}\sum_{t=0}^{T-1}D_{cham}(V^{\prime}_{t},V_{t})$, where $D_{cham}$ denotes the Chamber distance [65], $V^{\prime}_{t}$ and $V_{t}$ denote the mesh vertex coordinates for the reconstruction and the ground truth respectively:
@@
In the VAE, the distribution of the latent space for $z_{a}$ is encouraged to be close to a prior Gaussian distribution. The KL divergence is defined between the latent distribution and the Gaussian prior distribution. To control the trade-off between distribution fitting and diversity, we adopt the $\beta$-VAE formulation [64]. The KL loss $\mathcal{L}_{KL}$ is formulated as
$\mathcal{L}_{KL}=\beta\cdot KL(\mathcal{N}(\mu,\Sigma)\parallel\mathcal{N}(0,\mathbf{I}))$ (7)
which encourages the latent space $\mathcal{N}(\mu,\Sigma)$ to be close to the prior Gaussian distribution $\mathcal{N}(0,\mathbf{I})$.
@@
In terms of implementation, the mesh encoder $M_{\text{enc}}$ has three GCN layers and one fully connected (FC) layer. The mesh decoder $M_{\text{dec}}$ is composed of five FC layers. The Transformer encoder $T_{\text{enc}}$ and decoder $T_{\text{dec}}$ consist of two layers, four attention heads, a feed-forward size of 1,024, and a dropout rate of 0.1. The latent vector dimensions for the mesh and condition were set to 64 and 32, respectively. The model contains approximately 69.71 million parameters and was trained on an NVIDIA RTX A6000 GPU (48 GB) using the Adam optimiser with a fixed learning rate of $10^{-4}$ for $300$ epochs. Training was performed with a batch size of one cardiac mesh sequence, consisting of 50 time frames. The cardiac mesh at each time frame consists of 22,043 vertices and 43,840 faces. The weights $\beta$ and $\lambda_{s}$ in the loss function were empirically set to 0.01 and 1.
@@
To represent a cardiac mesh sequence in a low-dimensional latent space, we extract a latent vector after the Transformer encoder $T_{\text{enc}}$ but before the reparameterisation step. The latent vector is calculated as the mean of the latent vectors at the Transformer encoder output layer across 50 time frames. For calculating the latent delta, we quantify the deviation of the latent vector of the real heart to the latent vector of a group of synthetic hearts of the same attributes. Given conditions $c$, 100 samples of the latent variable $z_{a}$ are drawn from a standard Gaussian distribution, $z_{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $z_{a}$ denotes the latent space after reparameterisation in the VAE formulation. Each sample $z_{a}$ is concatenated with the condition latent vector $z_{c}$ and passed through the Transformer decoder and mesh decoder to generate a synthetic cardiac mesh sequence. Following synthetic mesh generation, each synthetic mesh sequence is provided to the mesh encoder $M_{\text{enc}}$ and Transformer encoder $T_{\text{enc}}$, to generate latent vectors across 50 time frames at the Transformer output later, subsequently averaged to form the latent vector $z^{synth}$. The real heart mesh sequence is provided to the mesh encoder $M_{\text{enc}}$ and Transformer encoder $T_{\text{enc}}$ for calculating the latent vector $z^{real}$ in the same manner.
With the latent vector $z^{real}$ for the real heart and the latent vector $z^{synth}$ for the synthetic heart, we define the latent vector as the Euclidean distance between $z^{real}$ and $z^{synth}$. As we draw 100 synthetic samples to represent a subpopulation with the same attributes, the latent delta $\Delta z$ is defined as:
$\Delta z=\left\|z^{\text{real}}-\frac{1}{100}\sum_{i=1}^{100}z^{\text{synth}}_{i}\right\|_{2}$ (10)
where $i$ denotes the sample index. The latent delta $\Delta z$ provides a robust metric to evaluate individual differences in cardiac structure and motion, quantifying the deviation of the real heart from its personalised normative model.
@@
This study used a data set of 38,309 participants obtained from the UK Biobank [46]. Each participant underwent cine cardiac magnetic resonance (CMR) imaging scans. From the cine CMR images, a 3D mesh sequence is derived to describe the shape and motion of the heart. The mesh sequence covers three anatomical structures, LV, Myo and RV. Each sequence contains 50 time frames over the course of a cardiac cycle. To derive cardiac meshes from the CMR images, automated segmentation [68] was applied to the images. The resulting segmentations were enhanced using an atlas-base approach [69], by registering multiple high-resolution cardiac atlases [70, 69] onto the segmentations followed by label fusion, resulting in high-resolution segmentations. A 3D template mesh [70] was then fitted to the high-resolution segmentations at the ED and ES frames using non-rigid image registration, generating ED and ES cardiac meshes. Subsequently, motion tracking was performed using Deepali [71], a GPU-accelerated version of the non-rigid registration toolbox MIRTK [72], on cardiac segmentations across the time frames. Deformation fields were derived using a free-form deformation (FFD) model with a control point spacing of $[8,8,8]$. The registration objective function included Dice similarity as the primary similarity metric and B-spline bending energy regularisation with a weight of 0.01. The deformation fields were derived between time frames and applied to propagate the ED mesh and ES mesh across the cardiac cycle. The proposed meshes were averaged using weighted interpolation based on temporal proximity to ED and ES [9] to ensure temporal smoothness of the resulting mesh sequence. All cardiac meshes maintained the same geometric structure.
The dataset was partitioned into training/validation/test sets for developing the MeshHeart model and a clinical analysis set for evaluating its performance for disease classification task. Briefly, MeshHeart was trained on 15,000 healthy subjects from the Cheadle imaging centre. For parameter tuning and performance evaluation, MeshHeart was evaluated on a validation set of 2,000 and a test set of 4,000 healthy subjects, from three different sites, Cheadle, Reading, and Newcastle centres. For clinical analysis, including performing the disease classification study and latent delta PheWAS, we used a separate set of 17,309 subjects from the three imaging centres, including 7,178 healthy subjects and 10,131 subjects with cardiac diseases and hypertension. PheWAS was undertaken using the PheWAS R package with clinical outcomes and coded phenotypes converted to 1,163 categorical PheCodes. P-values were deemed significant with Bonferroni adjustment for the number of PheCodes. The details of the dataset split and the definition of disease code are described in Supplementary Table S1.
@@
The code for this research is available on GitHub (https://github.com/MengyunQ/MeshHeart).